OpenStudy (anonymous):

Scientists estimate Earth's age to be 4.6 billion years. Samarium has a half-life of approximately 106 billion years. From the present date, approximately how many more years must pass before a sample of samarium that was created at the beginning of Earth's history contains 25 percent of the original parent material? 10 billion 26 billion 100 billion 206 billion

OpenStudy (anonymous):

\[\ln( N/N _{0})=-(\ln 2 \times t)/t _{1/2}\] \[N/N _{0}=0.25\] solving you get \[t=2 \times t _{1/2}\] Now subtract Earth's age \[t-4.6 \times 10^{9}\] you get the result you need.

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

what?